An arithmetic sequence has t4=4 and t10=5.5. What is the 88th term for the sequence? I said 25 but I only guessed correctly, I don't understand how to do it :(
Answer by scott8148(6628)

in an arithmetic sequence, the difference between consecutive terms is constant

there are 6 terms from t4 to t10, so the difference is (5.5-4)/6 or .25

there are 84 terms from t4 to t88, so t88=t4+(84(.25))___ t88=4+21